OpenStudy (anonymous):

How to do this problem? Must find the dy/dx y =log 10 e^x

OpenStudy (triciaal):

@Nnesha

OpenStudy (triciaal):

@DanJS

Nnesha (nnesha):

hmm sorry i havn't learned this stuff yet so sorry :(

OpenStudy (displayerror):

This is pretty helpful http://oregonstate.edu/instruct/mth251/cq/Stage6/Lesson/logDeriv.html If you don't want to look at the website, the general formula for the derivative of a log of some arbitrary base b is: \[\frac{d}{dx} \log_a(u)= \frac{1}{u \ln(a)} \frac{du}{dx}\] So with the parameters from your problem plugged in, you'd get this: \[\frac{d}{dx} \log_{10}(e^x)= \frac{1}{e^x \ln(10)} \frac{d\left(e^x\right)}{dx}\]